OpenStudy (anonymous):

Find X −3(x + 3) = −3(x + 1) − 5 Choices are: 6 -6 all real numbers no soulution

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

-3x-9=-3x-3-5 Adding both sides by +3x we get x's variable as 0; therfore no solution

OpenStudy (anonymous):

−3(x + 3) = −3(x + 1) − 5 -3x - 9 = -3x -3 -5 -3x - 9 = -3x -8 +3x +3x -9=-8 no solution